C3H8 + 5 O2 - - - - - - - - > 3 CO2 + 4 H2O
Calculate the mass of CO2 produced when 6.5 g of propane is reacted with 14.2 g of O2.
1
Expert's answer
2020-06-25T07:53:01-0400

First step is to determine the limiting reactant:

n (C3H8) = 6.5g / 44.1 g/mol = 0.147 mol

n (O2) = 14.2g / 32 g/mol = 0.444 mol

According to the equation, the C3H8 : O2 mole ratio is 1 : 5.

The actual ratio is 0.147 : 0.444 = 1 : 3 => O2 is the limiting reactant.

Now moles and mass of CO2 can be calculated.

n (CO2) = 3/5 x n(O2) = 0.267 mol

m (CO2) = 0.267 mol x 44 g/mol = 11.7 g

Answer: 11.7 g